Imagine you are running a new company that wants to start small and then scale. As the manager, you start out as the only full-time employee and you have a part-time support staffer and a part-time marketing employee.

1. At the beginning of your operations (year 1), the price of your product is $100. Your competitors are charging $110. Your direct costs are $60/unit ($40 labor, $20 materials). Your monthly overhead is as follows: facilities all-in $3,000, management $10,000, sales-marketing $6,000, support staff $2,500, IT-Telecom $500, accounting/legal/consulting $2,000.

What monthly volume of sales do you need to break even in year 1? ____________units/month

B) Now you are in the second year of operations. You have doubled your volume from year 1 where you exactly broke even. For year 3, you decide to expand the scale of your operation. You are hiring a second full-time manager and bringing on full time sales and marketing staff.

IIBIf you expect to double sales again from year 2 to 3 (thus quadrupling your year 1 volume), how much can you add to your staffing and management salaries and still break even? (Hint: Use the Burn section of Module C)

$______________/month

C) In year 4, competition is tightening in the market. Your competitors are now charging $100 and it is impacting your volume. With a burn rate of $48,000/month, monthly sales of 3000 units, and the same direct costs as before, what is the lowest price you could charge and still break even? (Hint: use the Pricing section of Module C)

$______________/unit

d) In year 5, you set your price at $90 and your burn rate is unchanged ($48,000). You expect to see 3600 units/month in sales. You decide to move to a more ecologically sustainable supply chain and you need to know what are the highest direct costs you can bear and still breakeven. What is the most you can pay in materials (assuming labor and other costs are unchanged)?

$______________/unit
Answer #1
Ans A-
Sales price per unit 100
Labor 40
Material 20 60
Contribution per unit 40
Overheads
facilities 3000
management 10000
sales marketing 6000
support staf 2500
IT telecom 500
accounting/legal 2000 24000
Break even sales per month 24000/40 =600
(total overheads/contribution per unit)
What monthly volume of sales do you need to break even in year 1? 600 units/month
Ans B- Production in second year =600*12*2= 14400
Production in third year =600*12*2*2= 28800
Contribution per unit 40
total Contribution (40*28800) 1152000
Monthly overheads 24000
yearly overheads (24000*12) 288000
Additional staff
second full time manager (12*10000) 120000
second full time sales staff (12*6000) 72000
Total overhead 480000
(288000+120000+72000)
Profit Per annum 672000
profit Per month 56000
To have still break even $ 56000 need to be added in staffing cost per month
Ans C- Monthly sales 3000 (per annum 3000*12) 36000 units
Direct cost (as before) 480000
Burn rate per annum 576000
(48000*12)
Net overhead after burn rate -96000
Variable cost per unit=60
total Variable cost (60*36000) 2160000
Net variable cost after burn rate 2064000
Net variable cost per unit after burn rate 57.33333
(=2064000/36000)
hence sales price can be set to the least @ $57.33 per unit to breakeven
Ans D- In year 5
sales price per unit $90
Units per year (3600*12)
sales (3600*12*90) 3888000
varable cost (60*3600*12) 2592000
Contribution (sales-variable cost) 1296000
Highest direct cost can be beared is $ 1296000 to have break even. 1296000
Hence per unit will be 1296000/432000= $3 per unit